The Conity Corporation has an Electric Mixer Division and an Electric Lamp Division. Of a $13,000,000 bond issuance, the Electric Mixer Division used $9,500,000 and the Electric Lamp Division used $3,500,000 for expansion. Interest costs on the bond totaled $975,000 for the year. What amount of interest costs should be allocated to the Electric Mixer Division? (Round any intermediary calculations two decimal places and your final answer to the nearest dollar.)

A) $563,333

B) $711,750

C) $3,500,000

D) $9,500,000
Answer #1

Answer : $ 711,750

The ratio of the Bond payable cost incurred between the two division

Electric Mixer Division = 9,500,000 / 13,000,000 = 73%

Electric Lamp division = 3,500,000 /13,000,000 = 27%

Interest Expesens Allocated to Electric Mixer Division = 975,000*73.% = $711,750 (Answer)